# Detection of Synthetic Face Images: Accuracy, Robustness, Generalization

### v1

@@ 4 Conclusion
We collected the FF5 dataset that includes images of recent diffusion generators. We conducted an extensive experimental study on the detection of synthetic face images. We will make our dataset and models publicly available. Our results allow us to draw the following conclusions.
@@
However, there are also bad news. It is simple to prepare an adversarial attack. It turns out that the residua found for a target model act adversarially, even on other models of very different architectures. We showed that adversarial images found for vision transformers often confuse convolutional networks. The worst news is that the detectors do not generalize well to generators they were not trained on. This is not just the case of our simple detector, but we showed that two tested state-of-the-art detectors failed completely and could not detect synthetic images generated by a newer generator, which they were not trained on.
These bad news demonstrate that universal detection of synthetic/fake images is a very tough problem, which is far from being solved despite active research in recent times.

### v2

@@ 4 Conclusion
In this paper, we conducted several experiments on the detection of synthetic face images. Our results allow us to draw the following conclusions.
@@
However, there are also bad news. It is simple to prepare an adversarial attack. It turns out that the residua found for a target model act adversarially, even on other models of very different architectures. We showed that adversarial images found for vision transformers often confuse convolutional networks. The worst news is that the detectors do not generalize well to generators they were not trained on. This is not just the case of our simple detector, but we showed that many tested state-of-the-art detectors could not reliably detect synthetic images generated by a newer generator, which they were not trained on.
This study presents multiple insights from targeted experiments and provides quantitative evidence on the challenges of synthetic image detection. In summary, current detectors trained discriminatively in a supervised manner learn to identify signal fingerprints of specific generators. Future research will focus on novel learning strategies to mitigate such overfitting and enhance generalization.
Acknowledgement. This work was supported by the national recovery plan projects CEDMO NPO (1.4 CEDMO 1 – Z220312000000) and CEDMO 2.0 NPO (MPO 60273/ 24/21300/21000) provided by the ministry of industry and trade, and by the CTU student grant SGS23/173/OHK3/3T/13. We thank to Patricie Petriľáková for evaluating the recent detectors on our dataset.